# A complex scalar field theory for charged fluids, superfluids, and fracton fluids

### v1

@@ 5 Fracton fluids
In this section, we explore this possibility and propose a fracton fluid phase, invariant under phase shifts linear in the comoving coordinates, corresponding to the comoving conservation of the dipole moment. An analogous fractonic solid phase was recently proposed by Akash Jain Jain:2024ngx  [^12] An idea to restrict the dynamics of charges with respect to the comoving observer was also discussed in Liang:2025gpp .. We begin by analysing the symmetry structure of fracton fluids and show that comoving multipole symmetries are generically incompatible with the full set of APDs, $\text{SDiff}(\mathbb{R}^{2})$, which must therefore be restricted to its affine subgroup, $\text{SL}(2,\mathbb{R})\ltimes\mathbb{R}^{2}$. Interestingly, the resulting symmetry group matches that of Ref. PhysRevD.89.045002 , which implemented a nonlinear realization of this group to approximate an ideal fluid.
Next, we formulate a symmetry-invariant action that is manifestly covariant under general coordinate transformations, local $\text{U}(1)_{Q}$ shifts, as well as local comoving dipole shifts. From this action, we derive the associated nonlinear hydrodynamic equations, construct a Goldstone EFT and compute the dispersion relations of the low-energy excitations. In addition to the ordinary sound mode, we identify a propagating “second sound” mode with a magnonlike dispersion, $\omega\sim k^{2}$. Hence, the physical content of the theory matches with the fractonic fluids described thus far in the literature Grosvenor:2021rrt ; Glorioso:2021bif ; Glodkowski:2022xje ; Armas:2023ouk ; Jain:2023nbf ; Glodkowski:2024ova ; Glorioso:2023chm , which conserve the dipole moment in physical space rather than in comoving space. In turn, the symmetry group implemented in PhysRevD.89.045002 is, at best, a crude approximation to ideal fluids, as it predicts additional gapless excitations that do not exist in a perfect fluid.
@@ 5.2 Relativistic model for fracton fluids
With these ingredients at hand, we are ready to formulate a field theory model for fracton fluids. Guided by Pretko:2018jbi ; Jain:2024ngx we propose the action
@@
We emphasize that the fracton fluid phase Eq. (114) is fundamentally different from the complex scalar field theories invariant under spacetime dipole symmetry Pretko:2018jbi ; Seiberg:2019vrp ; PhysRevResearch.2.023267 ; Bidussi:2021nmp ; Jain:2021ibh ; Jensen:2022ibn . In particular, the ordinary multipole symmetries, as classified by Gromov Gromov:2018nbv , are incompatible with Lorentz and Galilean boosts whereas Eq. (114) is exactly invariant under Poincaré transformations [^13] See also Bertolini:2022ijb ; Afxonidis:2023pdq for a Lorentz-covariant generalization of dipole symmetry based on the conservation of the four-dipole moment.. Moreover, the theory Eq. (114) is covariantly coupled to the background Lorentzian geometry through the spacetime metric $g_{\mu\nu}$. On the other hand, dipole-conserving theories are notoriously difficult to reconcile with gravity and instead couple to Aristotelian background geometries Pena-Benitez:2021ipo ; Pena-Benitez:2023aat ; Jain:2021ibh ; Grosvenor:2021hkn ; Hartong:2024hvs ; Afxonidis:2025wce .
The fracton fluid theory Eq. (114) is more closely related to the factonic solid phase introduced in Ref. Jain:2024ngx but differs in several key aspects. First, the fracton fluid is invariant under the $\text{SL}(2,\mathbb{R})\ltimes\mathbb{R}^{2}$ symmetry group, in contrast to the $\text{SO}(2)\ltimes\mathbb{R}^{2}$ symmetry that characterizes an isotropic solid or ”jelly” phase. This additional symmetry modifies the structure of the Lagrangian and, as a consequence, Eq. (114) does not support transverse phonon excitations. Second, the theory Eq. (114) is linear in time derivatives of $\Phi$ rather than quadratic. Therefore, the model Jain:2024ngx exhibits a Higgs mode in the broken phase whereas the fracton fluid theory does not. Finally, in our formulation, the comoving multipole symmetries arise naturally in the description of the normal phase, and the fracton fluid is introduced as an interpolating state between the normal and superfluid phases. On the other hand, the corresponding ”crystal-multipole symmetries” introduced in Ref. Jain:2024ngx are postulated from the outset.
@@ 5.4 Excitation spectrum
The spectrum of low-energy excitations is in qualitative agreement with previous studies on s-wave fracton superfluids with dipole moment conserved in physical space Armas:2023ouk ; Glodkowski:2024ova ; Jain:2023nbf ; Jain:2024kri . However, contrary to previous approaches utilizing the hydrodynamic paradigm, our derivation follows directly from the microscopic model Eq. (114). The fractonic two-fluid model Eq. (114), incorporating both the fractonic charge sector and the normal (thermal) component, generalizes the zero-temperature fractonic superfluid phases PhysRevResearch.2.023267 to a finite temperature regime.
The authors of Ref. PhysRevD.89.045002 employed the same symmetry group Eq. (109) and symmetry-breaking pattern Eq. (134) to model an ideal fluid as a nonlinear realization thereof. Our analysis of the microscopic model with the same symmetries shows, however, that this symmetry group predicts an additional gapless excitation with quadratic dispersion, which has no analogue in ordinary fluids. This indicates that the fractonic symmetry group Eq. (109) does not describe ordinary charge-carrying fluids, which must instead exhibit invariance under the full chemical shift symmetry, as discussed around Eq. (40).
@@ 6 Conclusion
In this work, we have presented a theory of charged fluids via a comoving hypersurface approach, with the complex matter field defined thereon. The proposed models, Eqs. (48) and (78), serve as a UV completion to the EFTs of charged fluids Dubovsky:2011sj and finite-temperature superfluids Nicolis:2011cs , respectively.
@@
Our framework provides a natural interpolation between normal fluids, whose charges are completely immobile on the comoving plane, and superfluids, whose charges are fully mobile. Such intermediate fluid phases are characterized by charges with partially restricted mobility on the comoving hypersurface, respecting a finite subgroup of chemical shifts corresponding to the comoving conservation of certain multipole moments. Focusing on the simplest case of fluids with comoving dipole symmetry, we have proposed a fracton fluid phase, in analogy with the fractonic solid phase introduced in Ref. Jain:2024ngx . We have shown that the physical content of the fracton fluid theory Eq. (114) accurately reflects the low-energy spectrum of fracton fluids studied extensively in the literature Armas:2023ouk ; Glodkowski:2024ova ; Jain:2023nbf ; Jain:2024kri , which respect dipole conservation in physical space.
@@
We conclude by outlining some interesting open problems for future exploration. Perhaps the most pressing question is the generalization of the comoving framework to include dissipative effects. Formulating dissipative hydrodynamics within an action principle allows for a systematic account of stochastic fluctuations, enables the computation of higher-point correlation functions, and ensures consistency with fluctuation–dissipation relations. While notable progress has been achieved Kovtun:2014hpa ; Crossley:2015evo ; Haehl:2018lcu ; Jain:2023obu ; Mullins:2025vqa ; Firat:2025upx , existing formulations are technically demanding and, in practice, difficult to use beyond reproducing established results, except in simple cases such as diffusive systems Delacretaz:2023ypv ; Delacretaz:2023pxm . In addition, such formulations typically rely on additional gauge symmetries whose physical provenance is not transparent. It would therefore be interesting to construct a dissipative EFT in terms of a complex scalar defined on the Schwinger–Keldysh contour over two copies of the comoving hypersurface.
@@
Another interesting direction is the potential adaptation of our construction to spinful fluids, which are of particular relevance for applications in heavy-ion collisions. Such systems are most naturally formulated in the tetrad formalism, coupled to the spin connection, possibly within a torsionful geometry Gallegos:2021bzp ; Gallegos:2022jow . To incorporate spin, one could introduce a set of fermionic fields, defined on the comoving hypersurface, in analogy with the bosonic matter fields discussed in this work. Notably, spinful fluids exhibit massive modes Hongo:2021ona , which may possibly be traced back to a more microscopic description within the comoving formalism.
@@
Finally, it would be interesting to modify our framework to incompressible quantum hall fluids and obtain a field-theoretical description of such phases, along the lines of Son:2013rqa ; Geracie:2014nka . For this purpose, one would need to replace the Poincaré symmetry used here with Galilean symmetry and place a system in a homogeneous background magnetic field. Subsequently, one could consider a formal lowest Landau level limit by sending the mass to zero $m\rightarrow 0$, which should imply the incompressibility constraint.
@@ Acknowledgements.
I am indebted to Piotr Surówka, Paweł Matus, Francisco Peña-Benítez, Sergej Moroz and Eren Fiat for useful discussions and comments. A.G. is supported in part by the Polish National Science Centre (NCN) Sonata Bis grant 2019/34/E/ST3/00405.
@@ Appendix B Representation of the comoving dipole symmetry group
This is precisely the algebra presented in the main text Eq. (107) albeit written in a different basis. Notice that (158) contains as a subalgebra a dipole algebra Gromov:2018nbv

### v2

@@ 5 Fracton fluids
In this section, we explore this possibility and propose a fracton fluid phase, invariant under phase shifts linear in the comoving coordinates, corresponding to the comoving conservation of the dipole moment. An analogous fractonic solid phase was recently proposed by Akash Jain Jain (2024) [^13] An idea to restrict the dynamics of charges with respect to the comoving observer was also discussed in Liang and Melia (2025).. We begin by analysing the symmetry structure of fracton fluids and show that comoving multipole symmetries are generically incompatible with the full set of APDs, $\text{SDiff}(\mathbb{R}^{2})$, which must therefore be restricted to its affine subgroup, $\text{SL}(2,\mathbb{R})\ltimes\mathbb{R}^{2}$. Interestingly, the resulting symmetry group matches that of Ref. Nicolis et al. (2014), which implemented a nonlinear realization of this group to approximate an ideal fluid.
Next, we formulate a symmetry-invariant action that is manifestly covariant under general coordinate transformations, local $\text{U}(1)_{Q}$ shifts, as well as local comoving dipole shifts. From this action, we derive the associated nonlinear hydrodynamic equations, construct a Goldstone EFT and compute the dispersion relations of the low-energy excitations. In addition to the ordinary sound mode, we identify a propagating “second sound” mode with a magnonlike dispersion, $\omega\sim k^{2}$. Hence, the physical content of the theory matches with the fractonic fluids described thus far in the literature Grosvenor et al. (2021); Glorioso et al. (2022); Głódkowski et al. (2023); Armas and Have (2024); Jain et al. (2023); Głódkowski et al. (2024); Glorioso et al. (2023), which conserve the dipole moment in physical space rather than in comoving space. In turn, the symmetry group implemented in Nicolis et al. (2014) is, at best, a crude approximation to ideal fluids, as it predicts additional gapless excitations that do not exist in a perfect fluid.
@@ 5.2 Relativistic model for fracton fluids
With these ingredients at hand, we are ready to formulate a field theory model for fracton fluids. Guided by Pretko (2018); Jain (2024) we propose the action
@@
We emphasize that the fracton fluid phase Eq. (114) is fundamentally different from the complex scalar field theories invariant under spacetime dipole symmetry Pretko (2018); Seiberg (2020); Yuan et al. (2020); Bidussi et al. (2022); Jain and Jensen (2022); Jensen and Raz (2024). In particular, the ordinary multipole symmetries, as classified by Gromov Gromov (2019), are incompatible with Lorentz and Galilean boosts whereas Eq. (114) is exactly invariant under Poincaré transformations [^14] See also Bertolini and Maggiore (2022); Afxonidis et al. (2024) for a Lorentz-covariant generalization of dipole symmetry based on the conservation of the four-dipole moment.. Moreover, the theory Eq. (114) is covariantly coupled to the background Lorentzian geometry through the spacetime metric $g_{\mu\nu}$. On the other hand, dipole-conserving theories are notoriously difficult to reconcile with gravity and instead couple to Aristotelian background geometries Peña-Benitez (2023); Peña-Benítez and Salgado-Rebolledo (2024); Jain and Jensen (2022); Grosvenor et al. (2022); Hartong et al. (2025); Afxonidis et al. (2025).
The fracton fluid theory Eq. (114) is more closely related to the factonic solid phase introduced in Ref. Jain (2024) but differs in several key aspects. First, the fracton fluid is invariant under the $\text{SL}(2,\mathbb{R})\ltimes\mathbb{R}^{2}$ symmetry group, in contrast to the $\text{SO}(2)\ltimes\mathbb{R}^{2}$ symmetry that characterizes an isotropic solid or ”jelly” phase. This additional symmetry modifies the structure of the Lagrangian and, as a consequence, Eq. (114) does not support transverse phonon excitations. Second, the theory Eq. (114) is linear in time derivatives of $\Phi$ rather than quadratic. Therefore, the model Jain (2024) exhibits a Higgs mode in the broken phase whereas the fracton fluid theory does not. Finally, in our formulation, the comoving multipole symmetries arise naturally in the description of the normal phase, and the fracton fluid is introduced as an interpolating state between the normal and superfluid phases. On the other hand, the corresponding ”crystal-multipole symmetries” introduced in Ref. Jain (2024) are postulated from the outset.
@@ 5.4 Excitation spectrum
The spectrum of low-energy excitations is in qualitative agreement with previous studies on s-wave fracton superfluids with dipole moment conserved in physical space Armas and Have (2024); Głódkowski et al. (2024); Jain et al. (2023, 2024). However, contrary to previous approaches utilizing the hydrodynamic paradigm, our derivation follows directly from the microscopic model Eq. (114). The fractonic two-fluid model Eq. (114), incorporating both the fractonic charge sector and the normal (thermal) component, generalizes the zero-temperature fractonic superfluid phases Yuan et al. (2020) to a finite temperature regime.
The authors of Ref. Nicolis et al. (2014) employed the same symmetry group Eq. (109) and symmetry-breaking pattern Eq. (134) to model an ideal fluid as a nonlinear realization thereof. Our analysis of the microscopic model with the same symmetries shows, however, that this symmetry group predicts an additional gapless excitation with quadratic dispersion, which has no analogue in ordinary fluids. This indicates that the fractonic symmetry group Eq. (109) does not describe ordinary charge-carrying fluids, which must instead exhibit invariance under the full chemical shift symmetry, as discussed around Eq. (40).
@@ 6 Conclusion
In this work, we have presented a theory of charged fluids via a comoving hypersurface approach, with the complex matter field defined thereon. The proposed models, Eqs. (48) and (78), serve as a UV completion to the EFTs of charged fluids Dubovsky et al. (2012) and finite-temperature superfluids Nicolis (2011), respectively.
@@
Our framework provides a natural interpolation between normal fluids, whose charges are completely immobile on the comoving plane, and superfluids, whose charges are fully mobile. Such intermediate fluid phases are characterized by charges with partially restricted mobility on the comoving hypersurface, respecting a finite subgroup of chemical shifts corresponding to the comoving conservation of certain multipole moments. Focusing on the simplest case of fluids with comoving dipole symmetry, we have proposed a fracton fluid phase, in analogy with the fractonic solid phase introduced in Ref. Jain (2024). We have shown that the physical content of the fracton fluid theory Eq. (114) accurately reflects the low-energy spectrum of fracton fluids studied extensively in the literature Armas and Have (2024); Głódkowski et al. (2024); Jain et al. (2023, 2024), which respect dipole conservation in physical space.
@@
We conclude by outlining some interesting open problems for future exploration. Perhaps the most pressing question is the generalization of the comoving framework to include dissipative effects. Formulating dissipative hydrodynamics within an action principle allows for a systematic account of stochastic fluctuations, enables the computation of higher-point correlation functions, and ensures consistency with fluctuation–dissipation relations. While notable progress has been achieved Kovtun et al. (2014); Crossley et al. (2017); Haehl et al. (2018); Jain and Kovtun (2024); Mullins et al. (2025); Firat et al. (2025), existing formulations are technically demanding and, in practice, difficult to use beyond reproducing established results, except in simple cases such as diffusive systems Delacretaz and Mishra (2024); Delacretaz (2025). In addition, such formulations typically rely on additional gauge symmetries whose physical provenance is not transparent. It would therefore be interesting to construct a dissipative EFT in terms of a complex scalar defined on the Schwinger–Keldysh contour over two copies of the comoving hypersurface. An open question is whether such a formulation reproduces all first-order dissipative transport coefficients, or whether some are missing due to the restricted field content of the comoving formulation.
@@
Another interesting direction is the potential adaptation of our construction to spinful fluids, which are of particular relevance for applications in heavy-ion collisions. Such systems are most naturally formulated in the tetrad formalism, coupled to the spin connection, possibly within a torsionful geometry Gallegos et al. (2021, 2023). To incorporate spin, one could introduce a set of fermionic fields, defined on the comoving hypersurface, in analogy with the bosonic matter fields discussed in this work. Notably, spinful fluids exhibit massive modes Hongo et al. (2021), which may possibly be traced back to a more microscopic description within the comoving formalism.
@@
Finally, it would be interesting to modify our framework to incompressible quantum hall fluids and obtain a field-theoretical description of such phases, along the lines of Son (2025); Geracie et al. (2015). For this purpose, one would need to replace the Poincaré symmetry used here with Galilean symmetry and place a system in a homogeneous background magnetic field. Subsequently, one could consider a formal lowest Landau level limit by sending the mass to zero $m\rightarrow 0$, which should imply the incompressibility constraint.
@@ Acknowledgements.
I am indebted to Piotr Surówka, Paweł Matus, Francisco Peña-Benítez, Sergej Moroz and Eren Firat for useful discussions and comments. A.G. is supported in part by the Polish National Science Centre (NCN) Sonata Bis grant 2019/34/E/ST3/00405.
@@ Appendix B Representation of the comoving dipole symmetry group
This is precisely the algebra presented in the main text Eq. (107) albeit written in a different basis. Notice that (158) contains as a subalgebra a dipole algebra Gromov (2019)